### v2

# Trustworthy Answers, Messier Data: Bridging the Gap in Low-Resource Retrieval-Augmented Generation for Domain Expert Systems

Abstract
RAG has become a key technique for enhancing LLMs by reducing hallucinations, especially in domain expert systems where LLMs may lack sufficient inherent knowledge. However, developing these systems in low-resource settings introduces several challenges: (1) handling heterogeneous data sources, (2) optimizing retrieval phase for trustworthy answers, and (3) evaluating generated answers across diverse aspects. To address these, we introduce a data generation pipeline that transforms raw multi-modal data into structured corpus and Q&A pairs, an advanced re-ranking phase improving retrieval precision, and a reference matching algorithm enhancing answer traceability. Applied to the automotive engineering domain, our system improves factual correctness (+1.94), informativeness (+1.16), and helpfulness (+1.67) over a non-RAG baseline, based on a 1-5 scale by an LLM judge. These results highlight the effectiveness of our approach across distinct aspects, with strong answer grounding and transparency.
Trustworthy Answers, Messier Data: Bridging the Gap in Low-Resource Retrieval-Augmented Generation for Domain Expert Systems
Nayoung Choi1, Grace Byun1, Andrew Chung1, Ellie S. Paek1, Shinsun Lee††thanks: Work done while at Emory University as a visiting scholar.1,2, Jinho D. Choi1 1Emory University, 2Hyundai Motor Company {nayoung.choi, grace.byun, andrew.chung, ellie.paek, shinsun.lee, jinho.choi}@emory.edu
1 Introduction
Retrieval-Augmented Generation (RAG) has shown potential in reducing hallucinations and providing up-to-date knowledge in Large Language Models (LLMs). This success has grown interest in domain expert RAG-Question Answering (QA) systems to meet specialized knowledge needs. While previous studies (Han et al. 2024; Siriwardhana et al. 2023; Mao et al. 2024) have proposed general methods for adapting RAG models to domain knowledge bases—such as syntactic QA pair generation or model fine-tuning—they face several challenges in low-resource domains.
In practical settings, available data sources in low-resource domains are often presented in heterogeneous formats and exhibit an unstructured nature, making their direct integration into RAG system development challenging Hong et al. (2024). General models may not have enough inherent knowledge in low-resource domains Zhao et al. (2024), making fine-tuning essential to adapt the model to specific knowledge requirements. However, the lack of structured data for training further complicates this process. In addition, data privacy and security concerns restrict the full utilization of API-based LLMs (Achiam et al. 2023; Anthropic 2024) within RAG systems, necessitating the use of open-source LLMs (Yang et al. 2024; Touvron et al. 2023).
The retrieval phase is another key aspect of domain expert RAG systems, as referencing accurate documents is essential for generating reliable answers. However, research on improving ranking in the retrieval phase or tracing the documents referenced to generate the answer remains limited. Most domain expert RAG frameworks rely on a single-stage retrieval process, with few exploring multi-stage approaches (Nogueira et al. 2019; Nogueira et al. 2020; Karpukhin et al. 2020), such as retrieval followed by re-ranking—a method widely used in Information Retrieval (IR)—which can help ensure the use of the most relevant references.
The evaluation of RAG systems is also an area that has not been fully addressed. Many studies continue to rely on gold answer similarity metrics, such as overlapping words between the generated answer and the ground truth, which inadequately capture critical dimensions like faithfulness, coherence, and contextual relevance. Recently, the LLM-as-a-judge framework Zheng et al. (2023) has gained attention as a qualitative alternative. However, these methods often overlook diverse evaluation aspects, and a standardized framework for assessing RAG systems has yet to emerge.
In this work, we address three key challenges and present the RAG development pipeline, demonstrating its application in the automotive engineering domain, with a specific focus on QA for vehicle crash collision tests. First, we present a data generation pipeline (Section 3.1), leveraging diverse formats of internal documents from an automobile company. Second, we incorporate an advanced re-ranking phase (Section 3.2) and a reference-matching algorithm (Section 3.4), not only enhancing the accuracy of the final answers but also improving traceability by tagging sources for each segment of the answer. Third, we evaluate the answers obtained from our generation model (Section 3.3) from multiple perspectives, emphasizing their qualitative aspects. Our ultimate goal is a fully local system, ensuring data privacy and independence from external servers. The overall workflow of our RAG system is shown in Figure 1, and our key contributions are summarized as follows:
We propose a data generation pipeline that transforms multi-modal raw data into a structured corpus and high-quality Q&A pairs.
We integrate re-ranking and reference matching to enhance retrieval precision and answer traceability, ensuring a reliable RAG system.111The full RAG pipeline code will be publicly available: https://github.com/emorynlp/MultimodalRAG
We assess the final answer from diverse qualitative perspectives, evaluating each along distinct, non-overlapping dimensions.
2 Related Work
Data Processing RAG-Studio (Mao et al. 2024) employs synthetic data generation for in-domain adaptation, reducing reliance on costly human-labeled datasets. However, it assumes access to well-structured data, limiting its applicability in scenarios with unstructured raw data. To bridge this gap, Hong et al. (2024) tackle challenges with real-world formats (e.g., DOC, HWP), proposing a chunking method that converts documents to HTML, retaining structural information in low-resource settings. Meanwhile, Guan et al. (2024) address the issue of short and noisy e-commerce datasets in RAG system development by building a new dataset from a raw corpus crawled from an e-commerce website, providing a richer resource.
Retrieval in RAG Wang et al. (2024) highlight the importance of re-ranking modules in RAG systems to enhance the relevance of the retrieved documents. Similarly, Zhao et al. (2024) demonstrate that the ranking position of the gold document in the retrieval phase plays a significant role in determining the quality of the final answer. Given these insights, optimizing the retrieval phase is crucial for obtaining accurate context, particularly in specialized, low-resource domains where LLMs lack sufficient inherent knowledge Beauchemin et al. (2024). Despite these findings, the analysis of the effectiveness and applicability of re-rankers in domain expert RAG systems remains underexplored.
RAG Evaluation The evaluation of RAG systems Yu et al. (2024) has relied on text similarity-based metrics such as BLEU Papineni et al. (2002), ROUGE Lin (2004), and EM Rajpurkar et al. (2016). These metrics provide a baseline but overlook valid answer diversity and qualitative aspects like factual consistency and relevance. Recent advancements are made to utilize LLM-as-a-judge Zheng et al. (2023) to evaluate qualitative dimensions. Han et al. (2024) propose a pairwise preference comparison framework considering helpfulness and truthfulness, while Saad-Falcon et al. (2024) assess context relevance, answer faithfulness, and query relevance, addressing hallucination issues. However, a standardized framework for RAG evaluation remains a challenge.
3 Approach
3.1 Data Generation & Processing
We build our dataset using two distinct sources: (1) internal reports on vehicle crash collision tests in presentation slide format from an automotive company in Korea, and (2) the textbook Crash Safety of Passenger Vehicles Mizuno (2016). Vehicle crash collision tests are exceptionally valuable, as each test incurs substantial costs and produces detailed reports that are typically confidential and difficult to access publicly. Figure 2 illustrates the process of extracting, analyzing, and converting slides and textbook PDFs into structured Markdown text and Q&A sets, leveraging Python scripts and the Claude LLM222We restrict the use of API-based LLMs for initial data processing; this limitation is discussed in Section 8. Anthropic (2024). Table 1 summarizes the data statistics by source.
The original slides often contain tables, graphs, and images, which are simplified into plain text descriptions during data processing. To evaluate performance on questions referencing these elements, we sample 143 multi-modal chunks and generated 1,861 targeted questions. Evaluation results are provided in Section 4.3.1.
Type | Source | File | Slide | Q&A Pairs |
PPT | Test Report | 1,463 | 4,662 | 59,402 |
Meeting Report | 249 | 882 | 7,696 | |
Page | Chapter | Q&A Pairs | ||
Textbook | 404 | 81 | 1,505 | |
Total | 5,625 | 68,603 |
Additionally, we extract and prepend headers summarizing each report, including the Test Name, Region, State, and Purpose, which significantly improve the retrieval phase (See Section 4.3). Domain experts are responsible for the entire process, including prompt engineering and reviewing intermediate and final outputs. All prompts and an example of a generated Q&A pair are shown in A.2.1.
3.2 Retrieval & Ranking
We employ a Dual-Encoder as our retriever, fine-tuning it on our training data. For a question , the model retrieves the top chunks from chunks , based on the [CLS] token embedding similarity between and each chunk , as follows:
Our re-ranker is a point-wise Cross-Encoder, trained on a classification task Nogueira et al. (2019) that takes a single pair as input and returns a scalar relevance score. It re-ranks , obtained from the retrieval stage, to extract , which will be passed to the generation model. The process is formalized as follows:
To optimize , we apply Token Selection (TS) + Term Control Layer (TCL) training method from RRADistill Choi et al. (2024). It effectively integrates the importance of the general semantic and specific query terms during the training process.
TS+TCL are used during training only, while inference uses the standard Cross-Encoder approach only with . The target label is binary: for relevant pairs and for irrelevant pairs. Negative sampling retrieves the top 10 chunks from the train data for each query , excluding the gold chunk, and randomly selects three from the rest.
3.3 Answer Generation
Our answer generation model takes the user question and the top chunks () from the retrieval and ranking phase as input to generate the final answer , as follow:
We fine-tune open-source LLMs on our training dataset, which consists of pairs derived from our data generation pipeline (Section 3.1). During fine-tuning, the question and the answer are concatenated into a single sequence , and the model is trained in an auto-regressive manner to predict the next token.
3.4 Reference Matching Algorithm
We propose a reference matching algorithm that segments generated answers and links them to relevant references using a Dynamic Programming (DP) approach Bellman (1954). Algorithm 1 outlines the detailed procedure.
In Step 1, the algorithm computes the scores for all possible sentence subsequences against the top-k chunks , using our re-ranker (Section 3.2). In Step 2, it selects the optimal segment-chunk combinations, updating scores and recording the best choices for each ending sentence (e.g., if the answer consists of five sentences like , the choices might be ). Finally, in Step 3, backtracking is performed to retrieve and output the best matches from the choices (e.g., ).
4 Experiment 1: Retrieval & Ranking
4.1 Models
For retriever training, we use BGE-M3 Chen et al. (2024) as the backbone, a multilingual Encoder capable of processing Korean, and fine-tune it on our training dataset with the publicly available code333https://github.com/FlagOpen/FlagEmbedding. For re-ranker training, we initialize the weights using the fine-tuned retriever. Further experimental details are in A.1.2.
4.2 Evaluation
We evaluate the retriever and re-ranker using Mean Average Precision (MAP@k) and Success@k Manning et al. (2008). MAP@k measures the ranking quality by averaging precision over relevant results up to rank , while Success@k indicates the proportion of queries with at least one relevant result in the top . The high Success@k score indicates that relevant chunks are retrieved, while the improved MAP@k highlights better ranking of those retrieved chunks. The evaluation is conducted using test set questions, but the chunk pool for retrieval included all splits (training, validation, and test) to ensure sufficient data and avoid biases from the limited test set size.
4.3 Result
Table 2 shows that prepending a header to each chunk—adding only a small number of tokens (see Table 9)—significantly enhances retrieval performance. Fine-tuning the BGE-M3 model on Ver.1 (BGE-FT) further improves results notably, demonstrating the importance of task-specific model adaptation to optimize performance. Table 3 shows that the re-ranking phase notably enhances the precision of retrieved information. The results indicate that ranking performance is optimal when the number of retrieved chunks is limited to 10. Despite the constraints of retriever failure, this improves retrieval results by about 4% in terms of MAP@1. In addition, incorporating TS+TCL during training boosts ranking performance (See A.3.1).
Model | Train | Test | MAP@1 | MAP@5 | MAP@10 |
BGE- Vanilla | N/A | Ver.0 | 0.1978 | 0.2672 | 0.2766 |
Ver.1 | 0.2991 | 0.3978 | 0.4091 | ||
BGE-FT | Ver.1 | Ver.1 | 0.6048 | 0.7111 | 0.7175 |
MAP@1 | MAP@5 | MAP@10 | Success@5 | |||||||||
10 |
20 |
30 |
4.3.1 Multi-modal Specific Questions
Table 4 summarizes the performance on questions requiring information from tables, graphs, and images. Compared to general questions (Section 4.3), the results are notably lower, with the retriever facing significant challenges. Even at = 30, Success@k only reaches 83.61%, underscoring a major bottleneck in the retrieval phase. Despite these challenges, the re-ranker proved effective, delivering a substantial performance improvement of approximately 15% in MAP@1 at = 20. These results underscore the need for further advancements to better handle multi-modal questions.
Phase | MAP@1 | MAP@5 | MAP@10 | |||
N/A | 0.3482 | 0.4561 | 0.4684 | ||
Reranking | 10 | 0.4723 | 0.5587 | 0.5630 | ||
20 | 0.4949 | 0.5923 | 0.5995 | |||
30 | 0.4933 | 0.5965 | 0.6041 |
4.4 Analysis
Figure 3 illustrates the trade-off between retrieval success and re-ranking performance as the number of retrieved chunks increases. While increasing improves the retrieval success rate (e.g., from 92% at = 10 to 96% at = 30), it can lead to diminishing improvements in re-ranking performance (e.g., MAP@5 decreases from 0.74 at = 10 to 0.73 at = 30). This result highlights the importance of carefully selecting to balance retrieval success and re-ranking quality, as overly increasing may not yield proportional benefits for re-ranking.
An analysis of the failed cases during the retrieval and ranking phase identified two error types: (a) Top-10 retrieval failure (23.3%) and (b) Retrieval success but incorrect re-ranker top-1 (76.7%). Type (a) were mainly caused by the open-ended nature of the questions, which led to multiple relevant documents beyond the gold context (Figure 18 in A.3.2). In contrast, type (b) occurred with more specific and detailed questions, where there were several relevant chunks from the same vehicle crash collision test, in addition to the gold chunk, making it difficult to identify a single correct one (Figure 19 in A.3.2). Failures in both types were often due to the presence of multiple valid chunks, rather than the ranking of irrelevant chunks. Table 4 presents a human evaluation of 100 sampled cases for each error type, assessing whether the re-ranker top-1 was relevant to the given question, even though it did not match the gold chunk. A screenshot of the human evaluation interface is in Figure 21 (A.5).
5 Experiment 2: Answer Generation
5.1 Models
We use Qwen-2.5 Yang et al. (2024) as our backbone LLMs, one of the few multilingual models that officially supports Korean. Notably, there are currently no Korean-centric open-source LLMs with sufficient context length to address our requirements. We fine-tune Qwen 2.5 (14B, 72B) on our training dataset, performing full fine-tuning on the 14B model and applying Low-Rank Adaptation (LoRA) (Hu et al., 2021) for the 72B model. Experimental details including hyperparameters and GPU configurations, are in A.1.2.
5.2 Evaluation
5.2.1 LLM-as-a-Judge
As illustrated in Figure 5, we employ the LLM-as-a-judge approach (Zheng et al., 2023) to evaluate the performance of both LLM-only and RAG models. For the LLM evaluation, we compare four model variations: the vanilla and fine-tuned versions of 14B and 72B models. GPT-4o444https://platform.openai.com/docs/models/gpt-4o is used to rank anonymized models (A, B, C, and D) based on factual correctness, helpfulness, and informativeness. These metrics are chosen to evaluate distinct, non-overlapping dimensions:
Correctness measures alignment with the gold answer, rewarding correct responses without hallucinations.
Helpfulness assesses clarity and relevance in addressing key points, while avoiding unnecessary content.
Informativeness evaluates the inclusion of relevant details or additional context that enhances completeness.
For RAG evaluation, we use the best-performing one among four models and compare its performance with and without RAG integration. Instead of pairwise comparisons, we employ single-answer evaluations, scoring responses (1–5) on three aspects. This approach offers a more detailed assessment, emphasizing how RAG impacts specific aspects of response quality. All the evaluation prompts used are detailed in A.2.2.
5.2.2 Human Evaluation
To assess the reliability of the RAG evaluation using the LLM-as-a-judge method, we conducted a human evaluation to compare its alignment with the LLM evaluation. An expert evaluator, a native Korean automotive engineer, assessed 100 randomly selected and anonymized responses. These responses were generated by the 72B fine-tuned model, both with and without RAG integration. A screenshot of the human evaluation interface used is in shown Figure 22 (A.5).
5.3 Result
5.3.1 LLM-only Comparison
Figure 6 illustrates the rank distribution and win-lose-tie comparison across four Qwen models. Among them, the 72B fine-tuned one shows the best performance, followed by the 72B vanilla, 14B fine-tuned, and 14B vanilla models. These results emphasize the critical impact of model size and demonstrate the significant effectiveness of fine-tuning, as fine-tuned models consistently outperform their vanilla counterparts.
5.3.2 LLM-only vs. RAG
We compare the performance of the 72B fine-tuned model in its LLM-only (non-RAG) and RAG-integrated versions. Table 5 compares the average scores for correctness, helpfulness, and informativeness between the two versions, while Figure 7 shows the win rates. The results clearly demonstrate that integrating RAG with the LLM improves factual correctness by reducing hallucinations and provides more helpful and informative answers.
Metric | non-RAG | RAG |
Correctness | 2.18 | 4.12 (+1.94) |
Helpfulness | 2.52 | 4.19 (+1.67) |
Informativeness | 2.61 | 3.77 (+1.16) |
5.3.3 Comparison with Human Evaluation
Table 6 shows that human evaluation aligns with GPT evaluation, both assigning higher scores to the RAG model across all three aspects. Figure 8 visualizes the differences in scoring tendencies between human and GPT evaluations. Notably, correctness and informativeness show an inverted pattern: human favors more polarized correctness scores (1 and 5), while GPT prefers mid-range (2 and 3). Additionally, human assigns higher helpfulness scores more frequently, as seen in the positive values for 4 and 5. While differences in score distributions exist between evaluators, these differences are independent of the specific models. Both evaluations consistently show that RAG outperforms non-RAG by achieving higher scores.
Metric | Model |
1 | 2 | 3 | 4 | 5 | ||
Correctness | A | 2.54 | 34 | 19 | 20 | 13 | 14 | ||
B | 4.33 | 10 | 4 | 4 | 7 | 75 | |||
Helpfulness | A | 2.89 | 2 | 38 | 36 | 17 | 7 | ||
B | 4.22 | 3 | 9 | 6 | 27 | 55 | |||
Informativeness | A | 2.60 | 3 | 44 | 44 | 8 | 1 | ||
B | 3.68 | 2 | 8 | 31 | 38 | 21 |
6 Experiment 3: Reference Matching
Question: Has there been an example where improving the vehicle structure prevented the fracture of components? | |||||
Segment 1 | Sentence 1 | Yes, there have been cases where improving the vehicle structure prevented the fracture of components. | Chunk 1,2,3 | ||
Segment 2 | Sentence 2 | First, in the ’HDC EV’ project, the design of the front bumper beam (FRT B/Beam) was changed to prevent fracture. | Chunk 2 | ||
Sentence 3 |
Segment 3 | Sentence 4 |
Chunk 1 | ||
Sentence 5 | For example, reinforcing the upper part of the A-pillar and the COWL BAR improved bending deformation. | ||||
Sentence 6 | Adding UPR HINGE B/H beads and other deformation-guiding methods reduced opening misalignment. | ||||
Segment 4 | Sentence 7 |
Chunk 3 | ||
Segment 5 | Sentence 8 | All of these improvements aim to optimize the body structure and prevent component fracture during collisions. | Chunk 1,2,3 |
6.1 Evaluation
To evaluate the reference matching task, we randomly sample 100 triplets () from the test set and perform human annotation to identify which chunk () each sentence in the generated answer () referenced. Each sentence is annotated with one or more chunks, allowing for multiple references. The interface of the annotation tool is in Figure 23 (A.5). We use sentence-level precision as the evaluation metric, defined as the proportion of sentences where the model prediction is included in the annotated reference set for an answer, averaged over the entire evaluation set. Since references are identified for concatenated sentences, selecting the best match for the grouped information is key. Recall would unfairly penalize valid predictions for not selecting every reference.
6.2 Result
Table 7 compares the reference matching performance of our algorithm (Section 3.4) without thresholding and the fine-tuned Qwen 72B, while Figure 10 illustrates the score distribution and the increase in precision with score thresholding for our algorithm. Without thresholding, our algorithm achieves a sentence-level precision of 0.72, while the LLM reaches 0.81, showing a noticeable gap in performance. However, when thresholding is applied, the precision reaches 0.86 at a threshold of 0.5, with performance improving proportionally as the threshold is increased further. This demonstrates the reliability of our algorithm’s scoring, allowing for controlled adjustment of reference matching quality through threshold selection. Our re-ranker used in the matching algorithm has only 0.5B parameters, compared to the LLM’s 72B, demonstrating that our algorithm can achieve impressive performance with a smaller model and faster inference speed. The LLM prompt used for reference matching is in Figure 17 (A.2.3). The answers are divided into 2-3 segments on average, and we identify two types of reference matching: one where all the necessary information to answer the question is contained within a single chunk, resulting in a single-segment answer, and another where information from multiple chunks is needed to answer the question, leading to a multi-segment answer. The first type often occurs in factual questions related to specific vehicle crash collision test, while the second type is more common in open-ended questions that require referencing multiple tests (Figure 9).
Method | Segment |
3.92 | ||||||
LLM |
13.54 |
7 Conclusion
In this work, we presented a low-resource approach to RAG for domain expert QA, utilizing heterogeneous data sources. We showed that our data generation pipeline, guided by expert-driven prompting, facilitates an effective RAG system by converting multi-modal raw data. By enhancing the retrieval phase with re-ranker, we demonstrated that the final answers are both accurate and traceable to relevant sources. These findings highlight the potential of RAG in data-scarce domains and its applicability to specialized knowledge-intensive tasks. As future work, we are exploring open-source Vision models to better handle multi-modal questions, addressing current limitations discussed in Section 8.
8 Limitations
Constraints in Backbone Model Selection The selection of a backbone model was constrained by the need for an open-source model supporting Korean with sufficient context length, limiting the available options across retrieval and LLM components. BGE-M3 for retriever & re-ranker, along with Qwen2.5 for the LLM, were the options available to us. However, it is worth noting that the newly released Qwen2.5-1M Yang et al. (2025) offers a longer context length, and our preliminary experiments demonstrated its potential for better handling long inputs. As the availability of open-source LLMs continues to grow, we anticipate that this limitation will gradually be resolved over time.
Restricted Use of API-based LLM for Initial Data Processing While our ultimate goal is to develop a fully local system, we initially relied on the API-based LLM, specifically the Claude API, for limited data processing. Among the available open-source LLMs, none met the requirements due to context limits and certain performance considerations, making them unsuitable for our needs. We determined that leveraging an API-based LLM with expert-driven prompting would be more effective in terms of both quality and cost compared to full human annotation. However, this reliance was limited to the initial data required for fine-tuning in the open-source model, and any future updates to the document pool will completely eliminate the need for external dependencies.
Weakness in Multi-modal Questions Our RAG system exhibited relatively lower performance when retrieving contexts for multi-modal specific questions that require information from tables, graphs, and images to answer the given question (Section 4.3.1). We hypothesize that this is due to the loss of rich content when multi-modal elements are converted into plain text during the data processing stage. To address this, we are experimenting with Llama 3.2 Vision555https://huggingface.co/meta-llama/Llama-3.2-11B-Vision to improve the conversion of multi-modal elements into text, aiming for better representation of these elements.
Experiments Limited by Computational Constraints The number of experiments we could conduct on the generation model was constrained by computational limitations. Fine-tuning the model on instead of (Section 3.3) proved impractical due to the extensive context length, which exponentially increased GPU memory requirements, surpassing our available computational resources. Although we observed that Qwen models inherently possess summarization capabilities for the given , and fine-tuning on yielded effective results, we were unable to explore a variety of learning strategies for the generation model within the RAG framework.
Appendix A Appendix
A.1 Experimental Details
A.1.1 Dataset
We split the generated dataset at the chunk level, slide level for slides, and chapter level for the textbook, using an 8:1:1 ratio for training, validation, and testing. Table 8 summarizes the dataset splits, while token length statistics for the training data are presented in Table 9.
Source | Data | Train | Val | Test | Total |
Test Report | Chunk | 3,729 | 466 | 467 | 4,662 |
Q&A Pair | 47,660 | 5,823 | 5,919 | 59,402 | |
Meeting Report | Chunk | 705 | 88 | 89 | 882 |
Q&A Pair | 6,144 | 752 | 800 | 7,696 | |
Textbook | Chunk | 64 | 8 | 9 | 81 |
Q&A Pair | 1,182 | 162 | 161 | 1,505 |
Min | Mean | Max | |
Question | 9 | 35 | 82 |
Answer | 6 | 56 | 271 |
Chunk (Ver.0) | 66 | 843 | 9,528 |
Chunk (Ver.1) | 106 | 884 | 9,528 |
A.1.2 Model Training
The retriever was trained for 10 epochs, determined to be optimal, using 4 48GB RTX A6000 GPUs. The hyperparameters were set to the default configuration provided in the open-source implementation of BGE-M3. For the ranker, early stopping was employed by validating the model every 1,000 steps. Training was stopped if performance did not improve after 3 consecutive validations, and the best-performing model was selected. This training used 1 48GB RTX A6000 GPU with a learning rate of 1e-5, the AdamW optimizer Loshchilov and Hutter (2017), and a batch size of 4. For both components, validation was based on the MAP@10 metric evaluated on the validation set. To train our answer generation model, full fine-tuning was conducted for the 14B model with a batch size of 2, gradient accumulation steps of 64, a learning rate of 2e-5, and 3 training epochs using 3 80GB H100 GPUs. For the 72B model, LoRA was applied with a batch size of 1, gradient accumulation steps of 64, a learning rate of 2e-5, and 3 training epochs using 2 × 141GB H200 GPUs.
A.2 Prompt Engineering
A.2.1 Prompts for Data Generation
The prompt used to convert slides into Markdown text is shown in Figure 12, and the prompt for Q&A generation is depicted in Figure 13. An example of a generated Q&A pair from a chunk is shown in Figure 14. An average of twelve Q&A pairs were generated from a single chunk.
A.2.2 Prompts for Final Answer Evaluation
Our prompts for evaluating final answers from LLM-only models and comparing LLM-only and RAG models are in Figure 15 and 16, respectively.
A.2.3 Prompts for Reference Matching
The prompt used for LLM reference matching is shown in Figure 17.
A.3 Further Analysis
A.3.1 TS+TCL Effect on Ranker Training
Table 10 shows the re-ranking performance without TS+TCL training. When compared to the performance with TS+TCL training (Table 3), there is a noticeable decline across all metrics.
MAP@1 | MAP@5 | MAP@10 | Success@5 | |||||||||
10 |
20 |
30 |
A.3.2 Analysis of Retrieval & Ranking Phase
Figure 18 and 19 present real-world examples of two error types: (a) cases where the top-retrieval process failed, and (b) cases where retrieval succeeded but the ranker’s top-1 result did not match the gold context.
Question: Were there cases where the body rating significantly decreased? | |||||||||||||||||||||||||||||||
Model Top-1 Chunk | Gold Chunk | ||||||||||||||||||||||||||||||
Model Top-1 Chunk | Gold Chunk | ||||||||||||||||||||||||||
A.3.3 LLM-only vs. RAG
Table 11 presents detailed counts and score distributions from GPT-4o, comparing RAG and non-RAG.
Metric | Score | Non-RAG | RAG |
Correctness | 1 | 1,640 | 388 |
2 | 3,166 | 579 | |
3 | 1,457 | 705 | |
4 | 430 | 1,359 | |
5 | 187 | 3,849 | |
Avg. Score | 2.1799 | 4.1195 | |
Helpfulness | 1 | 758 | 154 |
2 | 2,819 | 522 | |
3 | 2,422 | 868 | |
4 | 722 | 1,649 | |
5 | 159 | 3,687 | |
Avg. Score | 2.5211 | 4.1908 | |
Informativeness | 1 | 461 | 142 |
2 | 2,764 | 564 | |
3 | 2,789 | 1,609 | |
4 | 756 | 2,974 | |
5 | 110 | 1,591 | |
Avg. Score | 2.6061 | 3.7715 |
A.4 Analysis of Reference Matching
An example where only a single chunk needs to be referred to answer the question is in Figure 20.
Question: What was the material of the back beam used in the HDC P5 40kph slope 30 (RH) test? | ||||
Answer | Segment 1 | Sentence 1 | The material of the back beam used in the HDC P5 40kph slope 30 (RH) test was aluminum. | Chunk 1 |
Sentence 2 | This information can be found in the test conditions section. |
A.5 Human Evaluation Interface
The human evaluation interfaces for assessing the retrieval and re-ranking phase, as well as the reliability of the RAG evaluation with LLM, are shown in Figures 21 and 22, respectively. The human evaluation interface used for annotating the evaluation dataset to aseess the reference matching task is shown in Figure 23.